# Topology of dyonic AdS black holes with quasitopological electromagnetism in Einstein-Gauss-Bonnet gravity

### v2

@@ III Topology of critical points
In this section, we will investigate the categorization of critical points for black holes in all dimensions of space-time. By analyzing the metric of black holes in equation (5), we can deduce the thermodynamic function representing temperature without pressure
Figure 4: $d=6(\alpha=0.5,\beta=0.1,Q_{e}=1,Q_{m}=2)$: Up panel: The red arrows represent the vector field $n$ for dyonic AdS black holes with QTE in the EGB background. The colored dots are $(r,\theta)=(1.489048,\pi/2)$, $(r,\theta)=(2.706456,\pi/2)$ and $(r,\theta)=(1.873569,\pi/2)$ which represent the critical points $CP_{1}$, $CP_{2}$ and $CP_{3}$, respectively. Down panel:The deflection angle $\Omega(\vartheta)$ as a function of $\vartheta$ for contours $C_{1}$ (green curve), $C_{2}$ (brown curve) and $C_{3}$ (blue curve) and $C_{4}$ (black curve).
Figure 5: $d=6(\alpha=1,\beta=0.1,Q_{e}=1,Q_{m}=2)$: Up panel: The red arrows represent the vector field $n$ for dyonic AdS black holes with QTE in the EGB background. The colored dots are $(r,\theta)=(1.008427,\pi/2)$, $(r,\theta)=(3.609964,\pi/2)$ and $(r,\theta)=(3.284875,\pi/2)$ which represent the critical points $CP_{1}$, $CP_{2}$ and $CP_{3}$, respectively. Down panel: The deflection angle $\Omega(\vartheta)$ as a function of $\vartheta$ for contours $C_{1}$ (green curve), $C_{2}$ (brown curve) and $C_{3}$ (blue curve).
Figure 6: $d=7$: Up panel: The red arrows represent the vector field $n$ for dyonic AdS black holes with QTE in the EGB background. The blue dot are $(r,\theta)=(0.929668,\pi/2)$, which represent the critical points $CP_{1}$. Down panel: The deflection angle $\Omega(\vartheta)$ as a function of $\vartheta$ for contours $C_{1}$ (blue curve), $C_{2}$ (green curve).
@@
$\Psi$ $=\frac{1}{4\pi\sin\theta r_{+}\left(6(d-4)(d-3)\alpha+r_{+}^{2}\right)}$ (25)
@@
×$\left.\times 2(d-5)(d-4)\alpha+r_{+}^{2}\right)+r_{+}^{8-2d}$
×$\left.\left.\times\left(Q_{m}^{2}+\frac{r_{+}^{4d}Q_{e}^{2}}{\left(r_{+}^{2d}+8\beta\Gamma(d-1)^{2}r_{+}^{4}Q_{m}^{2}\right)^{2}}\right)\right)\right],$
@@
Next, we investigate the topological charge linked to the critical point across various dimensions. Through analysis and examination of the topological structure, the topological charge $Q$ is assigned a specific value for each critical point, when encircling a single critical point $CP$ with a contour $C$, it is observed that its non-zero topological charge $Q$ signifies significant information associated with the enclosed area. Conversely, if contour $C$ lies outside of critical point, its topological charge becomes zero cq6 ; cq7 . If there are many critical points within the contour, the total topological charge $Q$ is obtained by aggregating the individual topological charges associated with each critical point. In the orthogonal plane, we employ the method of Duan’s $\varphi$-mapping theory to ensure that the constructed contour exhibits a positive value. We can transform the critical point on the $(S,\theta)$ plane to the $(r,\theta)$ plane using reparameterization. In this case, the critical point is situated at $(r_{0},\frac{\pi}{2})$
@@
where $\vartheta\subset(0,2\pi)$. After this transformation, we can express the deflection angle
@@
Now, let us investigate the classification of critical points pertaining to AdS black holes. Research has indicated that black holes demonstrate characteristic swallowtail behavior across various dimensions of space-time and undergo small/large BH phase transitions resembling those observed in van der Waals liquid/gas systems cq3 . Interestingly, in the case of $d=6$, when the coupling constant $\alpha$ falls within a particular parameter range $(\alpha\geq 0.5)$, a noteworthy phenomenon occurs where the black hole experiences a phase transition characterized by small/intermediate/large BH phases. Additionally, this leads to the emergence of a triple points naturally, signifying that the black hole possesses three critical points. In the case where the coupling constant $\alpha=1$, it can be observed that despite the presence of three critical points, there are indications of the triple point phase transition within the system as suggested by the existence of two swallowtails. We aim to topologically classify the critical points for the dyonic AdS BHs with QTE in the EGB background. Based on the data presented in TABLE I, it is clear that when $d$ takes values of 4, 5, 6 and 7, there exists only one critical point for the BH, which undergoes a small/large phase transition in the system. For the case of $d=4$, the characteristics of $\theta$ and radius $r_{+}$ are presented in relation to the vector field $n$ in FIG. 1. Identify a critical point located at $(\theta,r_{+})$=$(\frac{\pi}{2},2.449490)$. Create two contours, namely $C_{1}$ and $C_{2}$, where $C_{1}$ should enclose critical point, while no critical points should be enclosed by $C_{2}$ (refer to TABLE II for contour parameters). In addition, the values of $\Omega(\vartheta)$ are observed to be $-2\pi$ and 00, respectively. The topological charge at the critical point $CP_{1}$ is $-1$. In contrast, contour $C_{2}$ did not enclose any critical point, resulting in a topological charge of zero. Inspired by cq4 , $CP_{1}$ is classified as a conventional critical point, and the total charge is $Q=Q_{CP_{1}}=-1$. The presence of only one critical point in the system, which is called as a physical critical point. Similarly, in the case of space-time dimensions $d=5$ and $7$, we observe analogous outcomes as for $d=4$. The distinction lies in the fact that the coordinates of the physical critical points differ $(d=5,(\theta,r_{+})=(\frac{\pi}{2},2.658528),d=7,(\theta,r_{+})=(\frac{\pi}{2},0.929668))$. Further details can be found in FIG.2 and FIG.6.
@@
$C_{1}$ $C_{2}$ $C_{3}$ $C_{4}$
$a$ 0.07 0.07 - -
($\alpha=0.5,d=4$) $b$ 0.4 0.4 - -
$r_{0}$ 2.449490 3.4 - -
$a$ 0.07 0.07 - -
($\alpha=0.5,d=5$) $b$ 0.4 0.4 - -
$r_{0}$ 2.658528 3.8 - -
$a$ 0.15 0.15 - -
($\alpha=0.1,d=6$) $b$ 0.4 0.4 - -
$r_{0}$ 2.57753 2.4 - -
$a$ 0.15 0.15 0.15 1.1
($\alpha=0.5,d=6$) $b$ 0.4 0.4 0.4 0.5
$r_{0}$ 1.489048 2.706456 1.873569 2.09775
$a$ 0.15 0.15 0.15 -
6($\alpha=1,d=6$) $b$ 0.4 0.4 0.4 -
$r_{0}$ 1.008427 3.609964 3.284875 -
$a$ 0.15 0.15 - -
6($\alpha=0.5,d=7$) $b$ 0.4 0.4 - -
$r_{0}$ 0.929668 2 - -
Next, let’s analyze the situation where $d=6.$
@@
The presence of a conventional critical point for a black hole can be observed at coordinates $(1.711584,\pi/2)$, with a topological charge of $(Q=Q_{CP_{1}}=-1)$ in FIG.3. This result exhibits remarkable similarities to findings in other dimensions $(d=4,5,7)$.
@@
We find that there are two swallowtails curves in the G-T diagram cq3 in the range of pressure $CP_{3}<P<CP_{2}$, we will discuss why the system has three critical points and only small/large BH phase transition occurs. The first observation is that the contours $C_{1}$ ,$C_{3}$ , and $C_{2}$ enclose the critical points $CP_{1}$ $(1.008427,\pi/2)$,$CP_{3}$ $(3.284875,\pi/2)$, $CP_{2}$ $(3.609964,\pi/2)$ respectively within the vector field in FIG.5. And the contours $C_{1}$ , $C_{2}$, and $C_{3}$ corresponding deflection angle $\Omega(\vartheta)=-2\pi,-2\pi$ and $2\pi$ for $\vartheta=2\pi$, respectively, this means that the black holes associated topological charge of $Q_{CP_{1}}=Q_{CP_{2}}=-$1, and $Q_{CP_{3}}=1$. As a result, the total topological charge can be determined by $Q=Q_{CP_{1}}+Q_{CP_{2}}+Q_{CP_{3}}=-1$. Naturally, the conventional critical point is denoted as $CP_{1}$ and $CP_{2}$, whereas the novel critical point is represented by $CP_{3}$. As the system undergoes small/large BH phase transition, it can be inferred that only $CP_{1}$ represents the conventional (physical) critical point. Furthermore, $CP_{2}$ and $CP_{3}$ are associated with non-physical critical points. In addition, the findings indicate that the critical point with topological charge of $-1$ can be considered as an indicator of the phase transition, whereas the novel critical point does not possess the ability to minimize Gibbs free energy and thus is defined as the critical point associated with non-physical characteristics ni1 . However, the conventional critical point $CP_{2}$ here lacks the ability to minimize the Gibbs free energy and is defined as a non-physical critical point. This suggests that the aforementioned conclusion cannot be universally applied to all systems.
Figure 7: a isobaric curves (colored solid curves) for dyonic AdS black holes with QTE in the EGB background. shown in the $T-r_{+}$ plane. The black curve is for the extremal points of the temperature. Up panel:$(\alpha=0.1)$,Middle panel:$(\alpha=0.5)$, Down panel:$(\alpha=1)$
@@ IV Nature of Critical points
In this section, we will investigate the nature of the critical points, it can be observed that the black hole experiences small/large phase transition with only one critical point $(d=6,\alpha=0.1)$. Similar conclusions can be drawn for $d=4$, $5$, and $7$. Hence, we will primarily focus on examining the critical point of the black hole at $d=6$. To this end, the equation of state (12) is employed to generate the isobaric curves(represented by colored curve) and the critical temperature (depicted as a black curve) in FIG.7. We can find some interesting results as follows:
@@
Up panel: Each isopiestic curve corresponds to distinct maximum extreme points of temperature, the red curve representing $P=0.02$, the green curve represents $P=0.025$, the blue curve means $P_{CP1}=0.027903$, and the orange curve indicates $P=0.03$. Additionally, it is noteworthy that the conventional critical point $(Q_{CP1}=-1)$ coincides with the maximum temperature extreme point. We present a $T-S$ equal-area diagram in FIG.8, here $T_{p}$ denotes the temperature at which the black hole experiences a phase transition.
Figure 8: $T-S$ diagram of the Maxwell ’s equal area law $(P=0.025,T_{p}=0.1349,\alpha=0.1,d=6)$.
The application of Maxwell’s equal area law, we can eliminate the unstable region near $CP_{1}$ (referred to as the intermediate black hole branch). Moreover, our findings reveal that the isobaric curves are divided into distinct regions with varying phase numbers by the critical temperature, and the arrangement of phases in proximity to the critical point is determined based on these regions. Interestingly, an increase in pressure $(red\rightarrow orange)$ results in a reduction of isopiestic lines near $CP_{1}$. Consequently, $CP_{1}$ can be regarded as a phase annihilation point.
@@
Middle panel: We can find that the isobaric curves $P_{CP3}=0.06353(purple)$; $P=0.006469(orange)$; $P_{CP1}=0.006536(blue)$; $P_{CP2}=0.0065757(red)$ exhibits temperature extreme points. We have observed that the maximum extreme temperature is associated with the green dots, which represent conventional critical points $CP_{1}$ and $CP_{2}$. On the other hand, the blue point (novel critical point $CP_{3}$) accurately represents the minimum extreme temperature. We can note that the region of $T-S$ is depicted in FIG.9, with $T_{p1}$ and $T_{p2}$ denoting the temperature of the black hole during small to intermediate and intermediate to large phase transitions, respectively. By applying Maxwell’s equal area law, we can effectively eliminate unstable regions near $CP_{1}$ and $CP_{2}$; however, this approach may not be successful for unstable regions near $CP_{3}$. These findings align perfectly with the results cq4 . Furthermore, as pressure increases (indicated by orange-red), we notice a decrease in the number of regions surrounding $CP_{1}$ and $CP_{2}$, as well as a decrease in the number of regions near $CP_{3}$.
Figure 9: $T-S$ diagram of the Maxwell ’s equal area law $(P=0.006492,T_{p1}=0.064718,T_{p2}=0.0647425,\alpha=0.5,d=6)$.
Figure 10: $T-S$ diagram of the Maxwell ’s equal area law $(P=0.025,T_{p}=0.05328,\alpha=1,d=6)$.
We designate $CP_{3}$ as the phase generation point while referring to critical points $CP_{1}$ and $CP_{2}$ as phase annihilation points.

### v3

@@ III Topology of critical points
In this section, we will investigate the categorization of critical points for black holes. By analyzing the metric of black holes in equation (26), we can deduce the thermodynamic function representing temperature without pressure
Figure 3: $(d=6,\alpha=0.5,\beta=0.1,Q_{e}=1,Q_{m}=2)$: Up panel: The red arrows represent the vector field $n$ for dyonic AdS black holes with QTE in the EGB background. The colored dots are $(r,\theta)=(1.489048,\pi/2)$, $(r,\theta)=(2.706456,\pi/2)$ and $(r,\theta)=(1.873569,\pi/2)$ which represent the critical points $CP_{1}$, $CP_{2}$ and $CP_{3}$, respectively. Down panel:The deflection angle $\Omega(\vartheta)$ as a function of $\vartheta$ for contours $C_{1}$ (green curve), $C_{2}$ (brown curve) and $C_{3}$ (blue curve) and $C_{4}$ (black curve).
@@
$r_{c}$ 0.864544 - -
($\alpha=0.5,d=8$) $T_{c}$ 0.173732 - -
$P_{c}$ 1.386830 - -
$r_{c}$ 0.845743 - -
($\alpha=0.5,d=9$) $T_{c}$ 0.235677 - -
$P_{c}$ 3.856570 - -
Figure 4: $(d=6,\alpha=1,\beta=0.1,Q_{e}=1,Q_{m}=2)$: Up panel: The red arrows represent the vector field $n$ for dyonic AdS black holes with QTE in the EGB background. The colored dots are $(r,\theta)=(1.008427,\pi/2)$, $(r,\theta)=(3.609964,\pi/2)$ and $(r,\theta)=(3.284875,\pi/2)$ which represent the critical points $CP_{1}$, $CP_{2}$ and $CP_{3}$, respectively. Down panel: The deflection angle $\Omega(\vartheta)$ as a function of $\vartheta$ for contours $C_{1}$ (green curve), $C_{2}$ (brown curve) and $C_{3}$ (blue curve).
$\Phi$ $=\frac{1}{4\pi\sin\theta r_{+}\left(6(d-4)(d-3)\alpha+r_{+}^{2}\right)}$ (33)
@@
×$\left.\times 2(d-5)(d-4)\alpha+r_{+}^{2}\right)+r_{+}^{8-2d}$
×$\left.\left.\times\left(Q_{m}^{2}+\frac{r_{+}^{4d}Q_{e}^{2}}{\left(r_{+}^{2d}+8\beta\Gamma(d-1)^{2}r_{+}^{4}Q_{m}^{2}\right)^{2}}\right)\right)\right].$
@@
Next, we investigate the topological charge linked to the critical point. Through analysis and examination of the topological structure, the topological charge $Q$ is assigned a specific value for each critical point, when encircling a single critical point $CP$ with a contour $C$, it is observed that its non-zero topological charge $Q$ signifies significant information associated with the enclosed area. Conversely, if contour $C$ lies outside of critical point, its topological charge becomes zero cq6 ; cq7 . If there are many critical points within the contour, the total topological charge $Q$ is obtained by aggregating the individual topological charges associated with each critical point. In the orthogonal plane, we employ the method of Duan’s $\varphi$-mapping theory to ensure that the constructed contour exhibits a positive value. We can transform the critical point on the $(S,\theta)$ plane to the $(r,\theta)$ plane using reparameterization. In this case, the critical point is situated at $(r_{0},\frac{\pi}{2})$
@@
where $\vartheta\subset(0,2\pi)$. After this transformation cq4 , we can express the deflection angle
@@
Now, let us investigate the classification of critical points pertaining to AdS black holes. Research has indicated that black holes demonstrate characteristic swallowtail behavior across various dimensions of space-time and undergo small/large BH phase transitions resembling those observed in van der Waals liquid/gas systems cq3 . Interestingly, in the case of $d=6$, when the coupling constant $\alpha$ falls within a particular parameter range $(\alpha\geq 0.5)$, a noteworthy phenomenon occurs where the black hole experiences a phase transition characterized by small/intermediate/large BH phases. Additionally, this leads to the emergence of a triple points naturally, signifying that the black hole possesses three critical points. In the case where the coupling constant $\alpha=1$, it can be observed that despite the presence of three critical points, there are indications of the triple point phase transition within the system as suggested by the existence of two swallowtails.
@@
$C_{1}$ $C_{2}$ $C_{3}$ $C_{4}$ $a$ 0.07 0.07 - - ($\alpha=0.5,d=5$) $b$ 0.4 0.4 - - $r_{0}$ 2.658528 3.8 - - $a$ 0.15 0.15 - - ($\alpha=0.1,d=6$) $b$ 0.4 0.4 - - $r_{0}$ 2.57753 2.4 - - $a$ 0.15 0.15 0.15 1.1 ($\alpha=0.5,d=6$) $b$ 0.4 0.4 0.4 0.5 $r_{0}$ 1.489048 2.706456 1.873569 2.09775 $a$ 0.15 0.15 0.15 - ($\alpha=1.0,d=6$) $b$ 0.4 0.4 0.4 - $r_{0}$ 1.008427 3.609964 3.284875 -
We aim to topologically classify the critical points for the dyonic AdS BHs with QTE in the EGB background. By assigning special values to the parameters and performing numerical calculations, we give the critical points of different dimensions in TABLE I. Noted that for $d\neq 6$, there is only one critical point for the parameter to be assigned different values, which undergoes a small/large phase transition in the system. For the case of $d=5$, the characteristics of $\theta$ and radius $r_{+}$ are presented in relation to the vector field $n$ in FIG. 1. Identify a critical point located at $(\theta,r_{+})$=$(\frac{\pi}{2},2.658528)$. Create two contours, namely $C_{1}$ and $C_{2}$, where $C_{1}$ should enclose critical point, while no critical points should be enclosed by $C_{2}$ (refer to TABLE II for contour parameters). Similarly, the test found that $d=7,8,9$, and the same situation occurs.
Now, let ’s analyze the complex situation with triple point $(d=6)$.
@@
The presence of a conventional critical point for a black hole can be observed at coordinates $(1.711584,\pi/2)$, with a topological charge of $(Q=Q_{CP_{1}}=-1)$ in FIG.2, there exists only one critical point for the BH $(\alpha=0.1)$, which undergoes a small/large phase transition in the system.
@@
We find that there are two swallowtails curves in the G-T diagram cq3 in the range of pressure $CP_{3}<P<CP_{2}$, we will discuss why the system has three critical points and only small/large BH phase transition occurs. The first observation is that the contours $C_{1}$ ,$C_{3}$ , and $C_{2}$ enclose the critical points $CP_{1}$ $(1.008427,\pi/2)$,$CP_{3}$ $(3.284875,\pi/2)$, $CP_{2}$ $(3.609964,\pi/2)$ respectively within the vector field in FIG.4. And the contours $C_{1}$ , $C_{2}$, and $C_{3}$ corresponding deflection angle $\Omega(\vartheta)=-2\pi,-2\pi$ and $2\pi$ for $\vartheta=2\pi$, respectively, this means that the black holes associated topological charge of $Q_{CP_{1}}=Q_{CP_{2}}=-$1, and $Q_{CP_{3}}=1$. As a result, the total topological charge can be determined by $Q=Q_{CP_{1}}+Q_{CP_{2}}+Q_{CP_{3}}=-1$. Naturally, the conventional critical point is denoted as $CP_{1}$ and $CP_{2}$, whereas the novel critical point is represented by $CP_{3}$. As the system undergoes small/large BH phase transition, it can be inferred that only $CP_{1}$ represents the conventional (physical) critical point. Furthermore, $CP_{2}$ and $CP_{3}$ are associated with non-physical critical points. In addition, the findings indicate that the critical point with topological charge of $-1$ can be considered as an indicator of the first order phase transition, whereas the novel critical point does not possess the ability to minimize Gibbs free energy and thus is defined as the critical point associated with non-physical characteristics ni1 . However, the conventional critical point $CP_{2}$ here lacks the ability to minimize the Gibbs free energy and is defined as a non-physical critical point. This suggests that the aforementioned conclusion cannot be universally applied to all systems.
[…]
@@ IV Nature of Critical points
In this section, we will investigate the nature of the critical points. We can find some interesting results as follows:
@@
Up panel: In FIG.5, we can find each isopiestic curve corresponds to distinct maximum extreme points of temperature, additionally, it is noteworthy that the conventional critical point $(Q_{CP1}=-1)$ coincides with the maximum temperature extreme point. We present a $T-S$ equal-area diagram in FIG.6, here $T_{p}$ denotes the temperature at which the black hole experiences a phase transition. The application of Maxwell’s equal area law, we can eliminate the unstable region near $CP_{1}$ (referred to as the intermediate black hole branch). Moreover, our findings reveal that the isobaric curves are divided into distinct regions with varying phase numbers by the critical temperature, and the arrangement of phases in proximity to the critical point is determined based on these regions.
Figure 6: $T-S$ diagram of the Maxwell’s equal area law $(P=0.025,T_{p}=0.1349,\alpha=0.1)$.
Interestingly, an increase in pressure $(red\rightarrow orange)$ results in a reduction of isopiestic lines near $CP_{1}$. Consequently, $CP_{1}$ can be regarded as a phase annihilation point, noted that $d=5,7,8,9$ is the same as this case.
@@
Middle panel: We have observed that the maximum extreme temperature is associated with the green dots, which represent conventional critical points $CP_{1}$ and $CP_{2}$. On the other hand, the blue point (novel critical point $CP_{3}$) accurately represents the minimum extreme temperature.
Figure 7: $T-S$ diagram of the Maxwell’s equal area law $(P=0.006492,T_{p1}=0.064718,T_{p2}=0.0647425,\alpha=0.5)$.
Figure 8: $T-S$ diagram of the Maxwell’s equal area law $(P=0.025,T_{p}=0.05328,\alpha=1)$.
We can note that the region of $T-S$ is depicted in FIG.7, with $T_{p1}$ and $T_{p2}$ denoting the temperature of the black hole during small to intermediate and intermediate to large phase transitions, respectively. By applying Maxwell’s equal area law, we can effectively eliminate unstable regions near $CP_{1}$ and $CP_{2}$; however, this approach may not be successful for unstable regions near $CP_{3}$. These findings align perfectly with the results cq4 . Furthermore, as pressure increases (indicated by orange-red), we notice a decrease in the number of regions surrounding $CP_{1}$ and $CP_{2}$, as well as a decrease in the number of regions near $CP_{3}$. We designate $CP_{3}$ as the phase generation point while referring to critical points $CP_{1}$ and $CP_{2}$ as phase annihilation points.